OpenStudy (anonymous):

In one day, a machine creates a new product and fails 11 times but passes quality control five times. What is the percentage of failures of the product? Round to the nearest tenth.

OpenStudy (anonymous):

I just answered this question.

OpenStudy (anonymous):

I did not see an answer to it I ma sorry

OpenStudy (anonymous):

Total made = 11 + 5 = 16 FAilures = 11 Percentage of failures = 11/16 = ?

OpenStudy (anonymous):

ty sir